OpenStudy (anonymous):

Can someone explain the difference between using the sine ratio to solve for a missing angle in a right triangle versus using the cosecant ratio?

OpenStudy (ipwnbunnies):

\[\sin(x) = \frac{ opposite }{ hypotenuse }\] \[\csc(x) = \frac{ 1 }{ \sin(x) } = \frac{hypotenuse}{opposite}\]